OpenStudy (anonymous):

For the function, f(x)=2x^3+3x^2-12x+1 use the first and second derivatives of f(x) to find the following. (3 points each) a. Find the intervals on which f is increasing or decreasing. b. Find the local maximum and minimum values of f(x). c. Find the intervals on which f is concave up and concave down. d. Find any inflections points. e. Find the maximum value of this function on [−2,4].

OpenStudy (solomonzelman):

What did you find so far?

OpenStudy (anonymous):

just started but

OpenStudy (anonymous):

6x^2+6x-12

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=2x^3+3x^2-12x+1 }\) \(\large\color{black}{ \displaystyle f'(x)=6x^2+6x-12 }\)

OpenStudy (anonymous):

6(^2+x-2)

OpenStudy (solomonzelman):

yes, I know what you meant (try to avoid such typos tho, please)

OpenStudy (anonymous):

x=1, x=-2

OpenStudy (solomonzelman):

Those are the roots of the derivative?

OpenStudy (solomonzelman):

You need, though, to solve two cases; \(\large\color{black}{ \displaystyle 0<6x^2+6x-12 }\) (the solution to this will tell you where f increases (since the derivative at these values is positive)) \(\large\color{black}{ \displaystyle 0>6x^2+6x-12 }\) (the solution to this will tell you where f decreases (since the derivative at these values is negative))

OpenStudy (anonymous):

critical numbers

OpenStudy (solomonzelman):

But you just need to find the intervals at which the function is increasing, and intervals at which the function is decreasing.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

that is increasing, and decreasing would be: 0<6(x-1)(x+2) x \(\in\) (-∞,-2) \(\cup\) (1,∞) 0>6(x-1)(x+2) x \(\in\) (-2, 1)

OpenStudy (solomonzelman):

Concave up, f''(x)>0 Concave down, f''(x)<0

OpenStudy (solomonzelman):

Points of inflection f''(x)=0

OpenStudy (solomonzelman):

e) - Find critical numbers (they are: -2,1,4) - Find which one is larger, f(-2), f(1) or f(4) ?

OpenStudy (anonymous):

i got everything exept c and d

OpenStudy (anonymous):

i don't understand them